Operation Summary for August 31, 2009 - September 2, 2009
Monday August 31
During the day shift, the water group replaced a temperature sensor at MI-30 after numerous LCW pump trips. Operations received permission for beam in the Pelletron, the Main Injector (MI), the 8 GeV Transfer Line, and MiniBooNE. The MI beam pipe repairs continued. Booster personnel investigated a kicker (MP02) regulation problem.
Tuesday Sept 1
During the midnight shift, the LRF2 vacuum pump tripped off and then, when operators reset it, tripped off the rack power. The pump power supply needs to be replaced.
During the evening shift, Pbar personnel investigated the vacuum problem. Booster tuning continued late into the shift. A Booster kicker (MP03) tripped off; it will get checked out in the morning. Operations monitored the Pelletron high voltage throughout the shift; there was no beam. Interlock technicians conducted safety system tests in MI.
Wednesday Sept 2
The Plans for Wednesday
FESS personnel reported that CUB Chiller #3 passed its nitrogen pressure test and its repair/installation continues. RF personnel reported that the Tevatron RF system was ready to go. The Switchyard EPB dipole magnets installation in the transfer hall should be completed today. Due to the Pbar vacuum problem, three sections of Pbar will have to be baked. MiniBooNE should receive beam today, after considerable tuning.
Comments and Suggestions
|last modified 09/04/2009 email Fermilab|
|Security, Privacy, Legal|